OpenStudy (anonymous):

A test of the prototype of a new automobile shows that the minimum distance for a controlled stop from 94 km/h to zero is 41 m. Find the acceleration, assuming it to be constant as a fraction of the free-fall acceleration. The acceleration due to gravity is 9.81 m/s 2 . Answer in units of g How much time does the car take to stop? Answer in units of s

OpenStudy (cwrw238):

use one of the equations of constant acceleration initial velocity u = 94 km / h, final velociy v = 0 distance s = 41 m, a = ? v^2 = u^2 + 2as plug in the values for u, v and s and solve for a oh - you'll need to convert 94 km / hour to m / s then a will be in m/s

OpenStudy (cwrw238):

* a will be in m s^-2 divide this by 9.81 to convert to units of g

OpenStudy (anonymous):

alright

OpenStudy (cwrw238):

94 km/h = 94,000 m / h = 94,000 / 3600 = 26.11 m s

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

how do you find a

OpenStudy (cwrw238):

so plugging in these vales into the formula: 0 = 26.11^2 + 2a*41 a = -(26.11^2 / 84 negative because the car is slowing down = -8 .116 m/s_ in terms of g this is -8.116 / 9.81 = -0.827 g

OpenStudy (cwrw238):

to find time to stop use v = u + at 0 = 26.11 - 8.116 t t = 26.11 / 8.116 = 3.22 seconds

OpenStudy (cwrw238):

do u follow that? have u got any questions?

OpenStudy (anonymous):

no thank you so very much

OpenStudy (cwrw238):

yw